Polluted water is passed through a series of filters. Each filter removes 59% of the remaining impurities. Initially the water contains impurities at a level of 400 parts per million (ppm). Determine a rule for the function g , that gives the remaining level of impurities, L , after the water has passed through a series of n filters.
G(n)=?????
Thanks!

Answer by Theo(13342)

each filter removes 59% of the remaining impurities.

remaining level of impurity equals 41% of the original impurity.

example:

400 parts per million start with.
pass through filter once.
remove .59 * 400 = 235
remaining = .41 * 400 = 164
235 + 164 = 400.

you can use .41 as the factor for the remaining impurities.

400 * .41 = 164 * .41 = 67.24 * .41 = 27.5684

basically you have 400 * .41 * .41 * .41 which is the same as 400 * .41^3

400 * .41^3 = 27.5684.

this works for any number of times you pass though the filter.

g(n) = 400 * .41^n

when n = 3, g(3) = 400 * .41^3 = 27.5684.

you could have used any variable name, such as x, as well.

g(x) = 400 * .41^x

when x = 3, g(3) = 400 * .41^3 = 27.5684.